OpenStudy (meilendurcer):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4. A. y = 3x - 12 B. y = -3x C. y = -3x + 12 D. y = 3x I think it's the third answer.

OpenStudy (kendricklamar2014):

its C

OpenStudy (meilendurcer):

@KendrickLamar2014 great thanks! can you possibly help with another?

OpenStudy (kendricklamar2014):

sure :D

OpenStudy (meilendurcer):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of one half. y - 4 = one half(x - 1) y + 4 = 2(x - 1) y + 4 = one half(x - 1) y - 4 = 2(x + 1)

OpenStudy (meilendurcer):

i'm not quite sure what the answer would be but my est guess is A

OpenStudy (kendricklamar2014):

Dont know how to do this one, sorry :(

OpenStudy (meilendurcer):

that's alright

OpenStudy (kendricklamar2014):

Sorry :(